A material has a mass of 83 g before abrasion and the abrasive ratio is 1.75. If the area of


abrasion is 23, find the mass after abrasion.

1
Expert's answer
2022-02-10T14:20:01-0500

Abrasion ratio= "\frac{175}{100}"


Mass before abrasion= 83g


Mass after abrasion= "\frac{100}{175}\times" 83g


Mass after abrasion= 47.429g